### v2

# Bose–Fermi $N$-polaron state emergence from correlation-mediated blocking of phase separation

Abstract
We study $N$ fermionic impurities in a one-dimensional lattice bosonic bath at unit filling. Using DMRG and mixed boundary conditions—closed for bosons, open for fermions— we find an N-polaron ground state replacing phase separation at high interspecies repulsion. This tightly bound state of clustered particles emerges due to strong impurity-bath correlations which induce large impurity-impurity correlations, that we quantify via the von Neumann entropy and bipartite mutual information respectively. This system also reveals a fermionic self-localization effect from a Mott insulator background due to local correlations between the impurities and the bath. The growth of long-range correlations breaks this Mott phase, resulting in the transition to localized impurity clusters. Finally, we show that there is a critical impurity number, which depends on intraspecies bosonic interaction, beyond which phase separation is recovered.
Since the seminal work by Landau, impurities inside their surrounding many-body media can be understood as entities wearing a “dress” of background excitation, leading to a notion of quasi-particles known as polarons [1, 2, 3]. Studying the properties of polarons is often the first step to tackling more complicated many-body systems [4, 5], providing further insight that could lead to theoretical developments in condensed matter systems, such as the mechanism of the high-temperature superconductivity [6, 7, 8, 9, 10, 11], one of the ultimate goals of modern many-body physics. As exemplified by this system, the presence of strong correlations among particles may lead to various phenomena that are fundamentally and practically important. The realization of polarons in cold atoms [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], opens a new and ideal platform to study correlation effects due to the high degree of tunability and clean environment [22, 23, 24, 25, 26, 27, 28, 29, 30]. This achievement has led to the study of the strong impurity-bath coupling regime, i.e. strongly correlated polarons, previously unattainable or uncontrollable in solid-state systems [31, 32, 33, 34, 35, 36, 37, 38, 39, 40], along with the investigation of the interplay between different particle statistics such as the Bose or Fermi polaron, an impurity immersed in a bosonic or fermionic bath [41, 42, 43, 44, 45], respectively, and multiple impurities where statistics and many-body effects play an important role [46, 47, 48, 49, 50, 51, 52].
In this work, we study a one-dimensional lattice system that consists of a few fermionic impurities immersed in a bosonic bath at unit-filling. This system falls under the umbrella of Bose–Fermi mixtures, for which phase separation is known to occur for strong interspecies repulsion [53, 54, 55, 56] and which is also true for some few-impurity systems [57, 46, 58]. In the latter, boundary and finite size effects significantly affect the ground state of the full many-body system which is the focus of our work. We elucidate this by comparing open boundary conditions (BCs) for both bosonic and fermionic systems (Open(B)-Open(F) BCs) with closed boundary conditions for bosons while keeping the fermionic boundary open (Closed(B)-Open(F) BCs). While the miscible-immiscible phase transition is observed as expected for Open(B)-Open(F) BCs, for high population imbalance the Closed(B)-Open(F) BCs system avoids phase separation at large interspecies interactions and instead forms a $N$-polaron state bounded by large bath-impurity correlations. We characterize this state using density profiles and appropriate correlation functions, along with tools developed in quantum information theory such as the Von Neumann entropy and mutual information, which are used extensively in the characterization of quantum phase transitions [59, 60, 61, 62, 63].
We consider a mixture of bosons and spin-polarized fermions loaded in a one-dimensional optical lattice of length $L$. The lattice depth is large enough that the low energy properties of the system can be well-described by a Bose–Fermi Hubbard model [64] with equal hopping parameters for the bosonic and fermionic species [65, 66]
$\hat{\mathcal{H}}$ =^$=\hat{\mathcal{H}}_{\rm B}+\hat{\mathcal{H}}_{\rm F}+\hat{\mathcal{H}}_{\rm BF},$ (1)
^$\hat{\mathcal{H}}_{\rm B}$ $=-t\sum_{j}\left(\hat{b}^{\dagger}_{j}\hat{b}_{j+1}+\rm{h.c}\right)+\frac{U_{\rm BB}}{2}\sum_{j}\hat{b}^{\dagger}_{j}\hat{b}^{\dagger}_{j}\hat{b}_{j}\hat{b}_{j},$ (2)
^$\hat{\mathcal{H}}_{\rm F}$ $=-t\sum_{j}\left(\hat{f}^{\dagger}_{j}\hat{f}_{j+1}+\rm{h.c}\right),$ (3)
^$\hat{\mathcal{H}}_{\rm BF}$ $=U_{\rm BF}\sum_{j}\hat{b}^{\dagger}_{j}\hat{f}^{\dagger}_{j}\hat{f}_{j}\hat{b}_{j},$ (4)
where $\hat{b}_{j}$ $(\hat{b}^{\dagger}_{j})$ and $\hat{f}_{j}$ $(\hat{f}^{\dagger}_{j})$ are the annihilation (creation) operators for bosons and fermions at the $j$-th site, respectively, with corresponding number operators $\hat{n}^{\rm B}_{j}=\hat{b}^{\dagger}_{j}\hat{b}_{j}$ and $\hat{n}^{\rm F}_{j}=\hat{f}^{\dagger}_{j}\hat{f}_{j}$. The terms $U_{\rm BB}$ and $U_{\rm BF}$ parameterize the strength of the boson-boson and boson-fermion coupling respectively, and will be varied to explore the phase diagram of the coupled system. In the following, we fix the number of bosons at unit-filling, $N_{\rm B}=L$, with system size $L=30$, while the number of fermions is much smaller than the background bosons, $N_{\rm F}\ll N_{\rm B}$. We initially fix $N_{\rm F}=2$ to illustrate the main physics and later discuss larger ratios of $N_{\rm F}/N_{\rm B}$. The ground state of the Hamiltonian (1) is found by using the Density Matrix Renormalization Group (DMRG) technique [67].
Figure 1: Local observables and bosonic energy gaps. Left and right rows correspond to Open(B)-Open(F) and Closed(B)-Open(F) boundary conditions, respectively. Panel (a) and (d): Density profile of the bosonic $\expectationvalue{\hat{n}^{\rm B}_{j}}$ and fermionic $\expectationvalue{\hat{n}^{\rm F}_{j}}$ species as a function of $U_{\rm{BF}}/U_{\rm{BB}}$ at $U_{\rm{BB}}=6$. Panel (b) and (e) show the normalized overlap $\Lambda/\Lambda_{0}$, and (e) and (f) the bosonic gap $\Delta_{\rm B}$ in logarithmic scale as a function of $U_{\rm{BF}}$ and $U_{\rm{BB}}$. For better visibility, we show the density cuts for the fermionic (blue circles) and bosonic (red squares) profiles at $U_{\rm{BF}}/U_{\rm{BB}}=0$ and $2$ in the upper panel of (a) and (d). Green and blue lines indicate the coupling strengths that give the maximum of the interspecies entropy $S_{\rm BF}$ (see Fig. 3) and fermionic relative distance $D_{\rm FF}$ [67], respectively. The dotted black lines indicate $U_{\rm BB}=U_{\rm BF}$ for reference.
We start by summarizing the case of Open(B)-Open(F) BCs and the well-known miscible-immiscible transition. In Fig. 1(a) we show the one-body density profiles of the mixture of the bosonic $\expectationvalue{\hat{n}^{\rm B}_{j}}$ and fermionic $\expectationvalue{\hat{n}^{\rm F}_{j}}$ species as a function of the interspecies coupling $U_{\rm BF}/U_{\rm BB}$ at fixed $U_{\rm BB}=6$. In the decoupled system ($U_{\rm{BF}}=0$, see left density cut from Fig. 1(a)), the background corresponds to a bosonic Mott insulator (MI) as the boson-boson interaction is much larger than the critical value of $U_{\rm{BB},0}^{\rm c}\approx 3.28$ [68, 69, 70, 71, 72] [^1] Although this critical point is measured in the thermodynamic limit, it is still an important reference value for the transition in this finite-sized system. , while the fermions are free and display two distinct peaks due to the Pauli repulsion. For sufficiently small interspecies interaction, $U_{\rm BF}\ll U_{\rm BB}$, the rigid insulating nature of the MI prevents any notable distortion to both density profiles; however, the density profiles change dramatically when $U_{\rm BF}\geq U_{\rm BB}$ as the system becomes phase separated with the fermions being individually localized at the box edges (see the right density cut in Fig. 1(a)).
The miscible-immiscible transition can be captured by the interspecies density overlap $\Lambda$ [74]
$\Lambda=\sum_{j}^{L}\expectationvalue{\hat{n}^{\rm B}_{j}}\expectationvalue{\hat{n}^{\rm F}_{j}},$ (5)
shown in Fig. 1(b) as a function of $U_{\rm BF}$ and $U_{\rm BB}$. We scale the overlap by the corresponding decoupled value $\Lambda_{0}(U_{\rm BB})\equiv\Lambda(U_{\rm BF}=0,U_{\rm BB})$ as a reference to the state of maximum miscibility for a given $U_{\rm BB}$. The critical value is clear at $U_{\rm BB}\approx U_{\rm BF}$, with an abrupt decrease in $\Lambda$ for larger interspecies couplings, $U_{\rm BF}>U_{\rm BB}$. The emergence of phase separation necessitates that the filling fraction of the bosonic background must increase above unity around the center of the box as the fermions are localized at the box edges, resulting in the loss of the Mott phase. This is signaled by the closing of the energy gap
$\Delta_{\rm B}=E^{L}_{N_{\rm B}+1}-2E^{L}_{N_{\rm B}}+E^{L}_{N_{\rm B}-1},$ (6)
where $E^{L}_{N_{\rm B}}$ is the ground state energy for $N_{\rm B}$ number of bosons with fixed $L$ number of lattice sites [75]. In Fig. 1(c), we see that the gap reduces for increasing $U_{\rm BF}$, saturating to vanishingly small values in the phase separation regime $U_{\rm BF}>U_{\rm BB}$, indicating that a superfluid (SF) background is necessary for the miscible-immiscible transition to occur.
The nature of the transition alters dramatically if one considers Closed(B)-Open(F) BCs, as seen from the density profiles in Fig. 1(d). In particular, the most relevant change occurs for $U_{\rm BF}\gtrapprox U_{\rm BB}$ where the absence of a boundary effect for the bosonic cloud prevents the system from transitioning to an immiscible state. Instead, the fermions reorganize themselves inside the bulk of the lattice on top of a reduced but relatively large bosonic accumulation (see right density cut from Fig. 1(d)). The absence of the immiscible state can be captured by $\Lambda$ (see Fig. 1(e)), where for fixed $U_{\rm BB}$ and increasing $U_{\rm BF}$, the overlap decreases smoothly rather than the abrupt transition in the Open(B)-Open(F) BCs case and saturates at a finite and relatively large value. Indeed, the density of the fermions does not change with increasing strong boson-fermion coupling $U_{BF}/U_{BB}\ggg 1$ (we have checked up to $U_{BF}/U_{BB}=20$, see Supp. Matt. [67]) with the fermions remaining spread over many lattice sites and any further localization being suppressed.
As is the case for the miscible-immiscible transition, the SF bosonic background is necessary for this re-ordering to occur, as indicated in Fig. 1(f), where $\Delta_{\rm B}$ behaves similarly to the Open(B)-Open(F) BCs case since this is a bulk property. Here, we note that the MI to SF transition is signaled by the maximum of the relative distance between the two fermions $D_{\rm FF}=\left(N_{\rm F}^{2}-N_{\rm F}\right)^{-1}\sum_{j,k}^{L}|k-j|\expectationvalue{\hat{f}^{\dagger}_{j}\hat{f}^{\dagger}_{k}\hat{f}_{k}\hat{f}_{j}}$ [57, 67] and is indicated by the blue lines in Fig. 1(d-f), marking the MI-SF transition of the bosonic system in $\Delta_{B}$ and bounding the region of maximum miscibility in $\Lambda$. In the MI phase, bosonic correlations are short-ranged [76], and therefore bosons only couple locally to the fermions, leading to weak self-localization of individual impurities as can be seen from Fig. 1(d). After the transition, long-range correlations that characterize the SF bath can, in turn, mediate long-range interactions between the impurities which leads them to localize together around the center of the lattice [67].
To further characterize the state of the fermions in the regime of strong impurity-bath coupling we calculate its one- and two-body reduced density matrices given by
$\hat{\rho}_{\rm F}^{(1)}$ $=$ $\frac{1}{N_{\rm F}}\expectationvalue{\hat{f}^{\dagger}_{j}\hat{f}_{k}},$ (7)
$\hat{\rho}_{\rm F}^{(2)}$ $=$ $\frac{1}{N_{\rm F}(N_{\rm F}-1)}\expectationvalue{\hat{f}^{\dagger}_{j}\hat{f}^{\dagger}_{m}\hat{f}_{l}\hat{f}_{k}},$ (8)
respectively. The one-body reduced density matrix is an auto-correlation function describing the probability of finding the same fermion simultaneously in two different lattice sites $j$ and $k$, while the two-body reduced density matrix describes the joint probability of finding one fermion in sites $j$ and $k$ and the other fermion in sites $l$ and $m$. As the latter is difficult to visualize we instead show in Fig. 2(a-c) its diagonal contribution, i.e. when $j=m$ and $k=l$, via the two-body density profile $\langle\hat{n}^{\rm{FF}}_{j,k}\rangle=(1-\delta_{j,k})\langle\hat{n}^{\rm F}_{j}\rangle\langle\hat{n}^{\rm F}_{k}\rangle$. This represents the probability of finding one fermion at site $j$ and the other at site $k$, and is, of course, zero at the diagonal ($j=k$) due to the Pauli exclusion principle. Free fermions are anti-correlated, as shown in Fig. 2(a), and the two particles are separated into two different sides of the box potential.
Figure 2: (a-c) Two-body fermionic density profiles $\langle\hat{n}^{\rm{FF}}_{j,k}\rangle$ and (d-f) one-body fermionic reduced density matrix $\hat{\rho}_{\rm F}^{(1)}$ for Closed(B)-Open(F) BCs. Here we have fixed the boson-boson coupling $U_{\rm BB}=6$ and take (a,d) $U_{\rm BF}=0$, (b,e) $U_{\rm BF}=1.333\,U_{\rm BB}$, and (c,f) $U_{\rm BF}=2\,U_{\rm BB}$.
However, for finite interspecies coupling $U_{\rm BF}=1.333U_{\rm BB}$ (Fig. 2(b)), for which the bosons are in the SF phase, the fermions start to bunch (they are more likely to be found near one another) and the relative distance between them decreases. For even larger $U_{\rm BF}$ (Fig. 2(c)) the fermions form a tightly bound state and $\langle\hat{n}^{\rm{FF}}_{j,k}\rangle$ is significantly broadened along the $j=k$ direction, indicating a large variance in the “center of mass” position and enhanced particle bunching (the fermions are most likely to be found in adjacent lattice sites). A similar behaviour can be seen in the one-body density matrix shown in Fig. 2(d-f), with the tightly bound state only displaying short-range off-diagonal coherences in Fig. 2(f). The emergence of this tightly bound fermionic state is due to effective attractive interactions that are mediated by the coupling to the bosonic background and are indicative of polaron physics [47, 50], showing that a bipolaron state is formed when considering Closed(B)-Open(F) BCs.
[…]
Figure 3: (a,c) Interspecies entropy $S_{\rm BF}$ and (b,d) fermionic mutual information $I_{\rm FF}$ as a function of $U_{\rm BF}/U_{\rm BB}$ for fixed $U_{\rm BB}=2$ (blue dot-dashed), $4$ (purple dashed) and $6$ (orange solid), and $I_{\rm FF}$ is normalized to the non-interacting value $I_{\rm{FF},0}=I_{\rm FF}(U_{\rm BF}=0)$. (a,b) correspond to Open(B)-Open(F) and (c-f) to Closed(B)-Open(F) BCs. The inset in (d) shows $U_{\rm BF}<U_{\rm BB}$ for $U_{\rm BB}=6$ with highlighted MI (gray) and SF (white) background phases, separated by the vertical solid blue line from Fig. 1(d). (e) $I_{\rm FF}$ vs $N_{\rm F}$ with $U_{\rm BF}\to\infty$ and (f) $I_{\rm FF}$ vs $U_{\rm BF}/U_{\rm BB}$ for $N_{\rm F}=15$ and the same $U_{\rm BB}$ values and labels as (a-d). The inset in (f) exhibits the fermionic (blue dots) and bosonic (red squares) density profiles for $U_{\rm BB}=6$ and $U_{\rm BF}=24$. Gray dotted lines indicate $U_{\rm BF}=U_{\rm BB}$ (vertical) and $I_{\rm FF}=I_{\rm{FF},0}$ (horizontal) for reference.
We first explore the case of Open(B)-Open(F) BCs where phase separation occurs, for which we show the von Neumann entropy and mutual information in Fig. 3(a) and (b) as a function of $U_{\rm BF}$. Starting from the decoupled pure states at $U_{\rm BF}=0$, $S_{\rm BF}$ increases with increasing $U_{\rm BF}$ as bosonic and fermionic states mix, reaching its maximum around the critical interaction strength $U_{\rm BF}\approx U_{\rm BB}$. In Fig. 1(a-c) we indicate the coupling strengths that give the maximum entropy with green lines, showing that it exactly separates the two phases and only shows very small deviations from $U_{\rm BF}=U_{\rm BB}$ below $U_{\rm BB,0}^{c}$. On the contrary, the mutual information $I_{\rm FF}$ decreases till the miscible-immiscible transition occurs, indicating impurity-impurity correlation screening as the impurities move to the system’s edges because of correlations with the bath, reaching a minimum near the $S_{\rm BF}$ maximum  [^2] For small $U_{\rm BB}$, the $S_{\rm BF}$ and $I_{\rm FF}$ extrema slightly differ but coincide as $U_{\rm BB}$ increases.. Increasing $U_{\rm BF}$ beyond this critical point leads to a decay in $S_{\rm BF}$ while $I_{\rm FF}$ returns to its initial value. These are both indicators of the phase separation regime in which the only contribution to the impurity-impurity correlations is the Pauli exclusion principle, while the bosonic bath acts as a double well potential that traps two impurities at the edges. The residual correlation in $S_{\rm BF}$ at high $U_{\rm BF}$ values is larger when the system is deeper in the SF regime ($U_{\rm BB}<U_{\rm BB,0}^{c}$), which can be understood as the compressible nature of the bosonic SF allowing a finite overlap with the fermions even at large $U_{\rm BF}$.
For Closed(B)-Open(F) BCs with weak interspecies coupling, $U_{\rm BF}<U_{\rm BB}$, $S_{\rm BF}$ and $I_{\rm FF}$ exhibit similar behavior as for Open(B)-Open(F) BCs, increasing and decreasing respectively until reaching a critical point that signals the miscible-polaronic transition. Again we take the coupling strength at the maximum of the von Neumann entropy and highlight it by the green lines in Fig. 1(d) and (e), which clearly signals the emergence of the bipolaron state. Moreover, in the inset of Fig. 3(d), we show in a closeup of the mutual information for $U_{\rm BB}=6$ that while in the MI phase (filled by grey background) $I_{\rm FF}$ is unchanged, indicating a lack of impurity-impurity correlations due to the “slow” spread of information through the insulating background [76], until the bath transitions to the SF state (white background). This transition almost coincides with the critical coupling strength obtained by the maximum relative distance between the impurities indicated by the blue vertical line, which is highlighted previously in Fig. 1(d-f). A similar behavior can also be seen in Fig. 3(b) for Open(B)-Open(F) BCs, although in this case, no clear indication of a transition appears in the fermionic density and distance like in the Closed(B)-Open(F) BCs case.
As expected from the differences in density profiles (Fig. 1(a,d)), the behavior changes radically after the miscible-polaron transition occurs where we observe a notable increase in $I_{\rm FF}$ and $S_{\rm BF}$ until both curves converge to large values in the $U_{\rm BF}\to\infty$ limit. Indeed, this indicates that strong impurity-bath correlations are responsible for the increased impurity-impurity correlations, as the former is necessary to induce large attractive interactions between the impurities and thus create the tightly bound states shown in Fig. 2(c). The saturation of the von Neumann entropy locks the correlations between the impurities and the bath, and therefore halts further localization of the fermions as shown in Fig. 1(d). This is an example of impurity self-localization suppression enhanced by impurity-bath correlations, an effect which was proposed in Ref. [78] using a variational ansatz in which the impurity-bath correlations are described nonperturbatively. Our work provides further evidence for this effect in discrete lattice systems, and here it is also responsible for the system to completely avoid the phase-separated state.
Finally, we note that the formation of $N$-polaron states through strong system-bath correlations is only present when there is a high interspecies population imbalance. To show this we analyze the regime of $N_{\rm F}>2$ where, practically speaking, obtaining $S_{\rm BF}$ is only possible for very few impurities since the dimension of $\hat{\rho}_{\rm F}$ increases as $L^{N_{\rm F}}\times L^{N_{\rm F}}$. In this case the bipartite mutual information $I_{\rm FF}$ represents the ideal tool as it signals the presence of bath-induced correlations and the formation of the $N$-polaron state when $I_{\rm FF}>I_{\rm FF,0}$, with $I_{\rm FF,0}$ being the mutual information of two fermions decoupled from the bosonic background ($U_{\rm BF}=0$). We show $I_{\rm FF}$ in Fig. 3(e) as a function of $N_{\rm F}$ and taking the limit $U_{\rm BF}\to\infty$ by removing the double Bose–Fermi occupancy in the Hilbert space for the DMRG calculations [67]. Adding impurities to the system for a fixed $U_{\rm BB}$ value effectively compresses the bath, as the bosons cannot occupy the same place as the fermionic cloud. This inhibits the bath-mediated interactions which reduces $I_{\rm FF}$ until a critical impurity density is reached and $I_{\rm FF}=I_{\rm{FF},0}$ indicating that the ground state is now phase separated. In Fig. 3(f), we show $I_{\rm FF}$ for the phase-separated state at fermionic half-filling $N_{\rm F}=15$, while the inset shows the density distributions at $U_{\rm BB}=6$ and $U_{\rm BF}=24$. In the phase-separated regime, the system exhibits a similar behavior as the Open(B)-Open(F) BCs case (Fig. 3(b)), however, the transition still occurs at larger values of the boson-fermion coupling $(U_{\rm BF}>U_{\rm BB})$. Finally, we note the crucial role that the bosonic interactions $U_{\rm BB}$ have on the critical impurity number. As can be seen from Fig. 3(e), for larger intraspecies interactions, phase separation occurs even for a small number of impurities as boson-boson correlations prevent overlap with the fermionic cloud, analogous to the behavior of $S_{\rm BF}$ at large $U_{\rm BF}$ in Fig. 3(a), therefore hastening the formation of the phase separated state.
We have studied the formation of a $N$-polaron state, where fermionic impurities are tightly bound together by a unit-filled bosonic gas in a one-dimensional optical lattice. We have shown that a superfluid-insulator transition is crucial for polaron formation, as when the Mott phase is broken and the background becomes a superfluid, correlations with the bosonic bath bind fermions together through strong mediated attractive interactions, allowing redistribution of the polaron across the superfluid bath and preventing phase separation, similar to Ref. [78] and recently observed in DMRG simulations [79]. This also suggests that the presence of strong impurity-bath correlations could inhibit the formation of the self-pinned state [80], as the mediated attractive interactions may prevent the impurities from self-organizing into a periodic structure. We have also shown that population imbalance is key for $N$-polaron formation, and that suitable tuning of bath intraspecies interactions and correlations could allow the creation of large and stable strongly-entangled $N$-polaron states. Indeed, a recent similar work has explored how the filling fraction and boson-boson interactions can affect bipolaron formation elucidating complementary results to our own [81]. Future studies could also investigate the role of particle statistics inside the polaronic state, analog to previous work with bosonic impurities using exact diagonalization [48, 51]. Moreover, the use of mixed boundary conditions on other well-known impurity systems could lead to novel phenomena by blocking or enhancing boundary effects.
Acknowledgements.
This work was supported by the Okinawa Institute of Science and Technology Graduate University. The authors are grateful for the Scientific Computing and Data Analysis (SCDA) section of the Research Support Division at OIST. The authors thank Tim Keller, L. A. Peña Ardila, G. A. Domínguez-Castro and F. Isaule for insightful discussions. T.F. acknowledges support from JSPS KAKENHI Grant No. JP23K03290. T.F. and T.B. are also supported by JST Grant No. JPMJPF2221.